Page tree
Skip to end of metadata
Go to start of metadata

 

Attendance: S. Wallace, E. Buchlin, T.Grundy, J. Lesho, S. Sidher, U. Schühle, D. Williams.

 

 

  1.     Highlights/recommendations
  • CMS operations to be performed weekly, capturing interesting events where possible. Interesting events since last meeting include the VGAM manoeuvres, whose results will be discussed in this month’s meeting 
  • CMS Operator should switch on CMS >30 minutes earlier than the time at which the observations would start, particularly if the observations to catch (for example) a WOL would need to start at midnight. CMS operations in SOOP Kitchen are timed such that they start 30 minutes after assumed switch-on. From LTP04 onwards, SOC will no longer indicate CMS operations at LTP with an RSW_EXT_SPICE. They will still be indicated with CMS “observations”, however.

 

 

  1.     CMS ops

In the past month, the spacecraft has had a superior conjunction with the Sun (Sun-Earth-SC angle < 5º) from end of 31st Jan until morning on 8th Feb. Bulk science data are still being caught-up from then (solid-state mass memory is in back-log). This should not affect HK data, via which CMS data come down (always brought down at start of the next available pass).

 

 

  1.     Anomalies (if any)

N/A                                           

 

 

  1.     CMS data and interpretation 
  2. a) See IAS CMS data tool

 


Operations times shown for CMS, since last Ops Team meeting


  1. STP-132: 2021-01-28T20:57 – 23:57 (DOY 028) around WOL
    • 22:42: seems to be some sort of unusal behaviour : 0.1 Hz change. Change infreq behaviour after 23:15. Small change.
    • Merits some investigation, since it doesn’t seem to be thruster firings.
    • Any manoevering? But it’s on the inside.
    • Maybe when SWA made some changes. Could be changes in the SPICE SEB temperature and the HIS bracket. AI: David Williams to check it out.
  2. STP-133: 2021-02-01T00:00 – 03:00 (DOY 032) around WOL – looks nominal
  3. STP-134: 2021-02-09T18:30 – 21:30 (DOY 040) around SA Relubrication – as expected
  4. STP-134: 2021-02-11T23:00 – 2021-02-12T02:00 (DOY 042–043) around 
    • HSD Opening for SPICE (and Metis) at 23:30:01, so CMS-1 expected to be relatively warm during the following 2 weeks : door can be detected opening at 23:30.
    • and WOL at 00:05 – 03:05
  5.             e. STP-136: 2021-02-22T00:00 – 03:00 (DOY 053) around WOL. – as expected.


Of interest to see how HSD being open for a couple of weeks affects measurements.


  1. Future operations:


Commanding for STP periods up to 137 (which ends 2021-03-08T00:00) have now been sent to MOC. 


Upcoming planned CMS Ops:

2021-02-28T09:30:00Z (SA Rotation) 60º – 70º

2021-03-08T20:30:00Z (TCM) ~6 hours measurement

2021-03-15T00:00:00Z (WOL)

2021-03-22T00:00:00Z (WOL)

 

Thermal measurements look very close to prediction.


 

  1.   AOB


On 9th Feb, had some bad data frames at end of the measurement. Occasional phenomenon, but the very last test-mode data frame normal but the samples were rubbish. Might occasionally get a byte-shift error, but this seems different: quite garbled. Means 1 sample less. CCSDS data are normal, but the data frame is affected. It HAS worked perfectly normally since then. This was warm, before the absolute maximum temperature was reached, before the HSD_OPEN.


SOC public states that MOC have distributed FOP 4.0. Seems ITs didn’t all get this. David Williams to alert SOC.


As May and June come up, so we may want to consider operations in that colder part of the orbit.



_______________________

 

Action Items

 

AI-12: MS to provide CMS1 & CMS2 FoV Open (DW will ask MS) important for next VGM

AI-13: David Williams to investigate 28th Jan 


_________________________


  • No labels